OpenStudy (anonymous):

g (h + 3/4)=11 M+F+T

OpenStudy (anonymous):

Solve for H

OpenStudy (pawanyadav):

You have to give us the function g and h For this

OpenStudy (anonymous):

?

OpenStudy (pawanyadav):

What is g...

OpenStudy (anonymous):

a letter?

OpenStudy (anonymous):

it just says solve for h using the literal equation x/16=a+g

OpenStudy (pawanyadav):

From this g=x/16-a Now what is a?

OpenStudy (anonymous):

im not sure..

OpenStudy (pawanyadav):

Then it should be a constant and you will get h in terms of a.

OpenStudy (anonymous):

so is that the answer

OpenStudy (pawanyadav):

No...

OpenStudy (anonymous):

can u tell me what it is?

OpenStudy (pawanyadav):

It should be like this g(x)=16 And g(x)=x/16 -a And now you have to put x=h+3/4

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

thats the answer right

OpenStudy (pawanyadav):

Welcome .very good if you got this...

OpenStudy (anonymous):

i didnt but thank u anyways

OpenStudy (pawanyadav):

Ha ha , I'll try to understand you..

OpenStudy (pawanyadav):

You are given a function g(x) in which if you put x=h+3/4 You will get 11.

OpenStudy (pawanyadav):

But there is a confusion as your question is not quite clear.... You are not given with the function g(x). g(x) must be an expression in x.

OpenStudy (pawanyadav):

Do you get it now...